# Rapid formation of a very massive star $>\!\!50000\,\mathrm{M}_{\odot}$ and subsequently an IMBH from runaway collisions

### v1

@@ A.1 Treatment of stellar winds for He star
The formation of VMS and posterior collapse as BH, is highly influenced by stellar winds, when it becomes an He star, which vary depending on the star metallicity. We updated the old treatment used in N-body for stellar wind from Hamann & Koesterke (1998); Vink & de Koter (2005) with the recipe from Sander & Vink (2020),
@@ Appendix B Upgrade of stellar radius evolution
Stellar evolution parameters, including the stellar radius, are provided by the SSE algorithm (Hurley et al. 2000a) with updates as described in Kamlah et al. (2022a). This algorithm was fitted to the results of detailed stellar models of up to 50.0 $\mathrm{M}_{\odot}$ and tested for reliable behaviour up to 100.0 $\mathrm{M}_{\odot}$. Thus, using SSE beyond this mass range was historically at the peril of the user and inconsistencies in the stellar radius for masses well above 100.0 $\mathrm{M}_{\odot}$ had been noted previously (e.g. Gieles, private correspondence). For the purpose of the Nbody6++GPU models reported in this work, we therefore conducted a series of tests of SSE in the VMS regime and implemented some minor additions to SSE to safeguard the routines when used for VMS evolution. In Appendix D, including Fig. 9, we explain in some more detail how the upgrade has been done. Note that this is aimed primarily at making the routines numerically safe to use as well as removing any obvious inconsistencies. In the long term, we plan to improve the code with routines for very low metallicities (Tanikawa et al. 2020) and massive star formation.
@@ Appendix C Upgrade of rejuvenation via stellar collisions
We use the BSE package by Hurley et al. (2002a, 2005), with upgrades related to the treatment of rejuvenation of main sequence stars upon collision with other MS stars and mass loss during hyperbolic collisions, as will be described in this Appendix. Rejuvenation originally was intended to describe the formation of blue straggler stars after stellar collisions (Hurley et al. 2005). These stars lie above a stellar populations main sequence turn-off point (see e.g., the observational evidence from Giesers et al. 2018). This observation implies that the effective age, $A_{\mathrm{MS}}$, of these MS stars after a collision and a merger, is younger than the average age of the stellar population. This concept (described in more detail below) delivers unphysical results for collisions of a VMS with small mass MS stars, i.e. those where the mass ratio $q$,
@@
becomes very small ($q\ll 1$). In the following, we will analyze in more detail the traditional rejuvenation treatment (Eq. 8, see also discussion in Hurley et al. 2005), explain why it fails for very small $q$, and describe our workaround (Eq. 10).
@@
    MS stars with convective cores $\left(M>1.25\leavevmode\nobreak\ \mathrm{M}_{\odot}\right)$: upon collision and subsequent mass gain by the primary MS star, its core grows in size. The latter leads to enhanced convective mixing of unburnt hydrogen fuel from the outer stellar layers to the core.This process will lead to the rejuvenation of the star. BSE numerically translate this rejuvenation by conserving the amount of burnt hydrogen and grown the stellar core of the primary star. The new age value is thus directly proportionally to the remaining fraction of unburnt hydrogen at the centre.
@@
    fully convective MS stars $\left(M<0.35\leavevmode\nobreak\ \mathrm{M}_{\odot}\right)$: these low-mass stars are treated in the same way as the stars with convective cores.
If we work under the base assumption that hydrogen in the MS core is uniformly distributed after a collision, which in turn is a result of the assumption of instantaneous (convective) mixing of the hydrogen fuel from envelope to core, the BSE code terminates the MS phase of the star and enters the Hertzsprung gap (HG) phase after 10 % of the total amount of hydrogen in the MS star have been burnt (Hurley et al. 2000a, 2005).
The computational application is described in the following. We first consider the collision of two MS stars with (effective) ages $A_{\mathrm{MS,1}}$ and $A_{\mathrm{MS,2}}$, MS life-times $\tau_{\mathrm{MS,1}}$ and $\tau_{\mathrm{MS,2}}$, as well as masses $M_{\mathrm{MS,1}}$ and $M_{\mathrm{MS,2}}$, respectively. Here, the subscripts $1$ and $2$ denote the primary (and thus more massive) MS star and the secondary, respectively. The collision produces the ”third” MS star with effective age $A_{\mathrm{MS,3}}$, MS life-time $\tau_{\mathrm{MS,3}}$ and mass $M_{\mathrm{MS,3}}$. If we know these parameters, (like suggested in Hurley et al. 2005) we can evaluate the final effective age $A_{\mathrm{H,MS,3}}$ of the produced MS star with the following equation (see also Glebbeek et al. 2008, for more information):
@@
Such nonphysical age reduction in the previous rejuvenation process has another unwelcome consequence, where the stellar radius of the star is always reset to a value that corresponds to the star’s effective age. However, a MS star grows in size as it ages and approaches the Hertzsprung gap (see e.g. Pols et al. 1998; Hurley et al. 2000a; Tanikawa et al. 2020). The stellar size increment accelerates the closer the star is located to the Hertzsprung gap. Therefore, the star will increase in size measured by the stellar radius in several orders of magnitude for VMSs. This implies that a MS star produced via repeated stellar collisions, whose effective age is treated with Eq. 8, will be much too small, which in turn greatly also affects its scattering cross-section and therefore, its mass growth, in dense stellar environment.
In our previous direct $N$-body works, where only a limited number of stellar collisions have been produced, no MS star merger chains have been produced that go beyond 10 compared with the densest star cluster model presented here with $r_{\mathrm{h,i}}=0.1\leavevmode\nobreak\ (\mathrm{pc})$ collisions (see e.g. Rizzuto et al. 2021, 2022; Kamlah et al. 2022a, b; Arca Sedda et al. 2023a, 2024b, 2024a, for recent direct $N$-body work, which includes comparable astrophysical complexity as the work presented here, i.e. full stellar evolution of single and binary stars, tidal fields, mass spectrum and so on.). Therefore the aforementioned rejuvenation issue was not a problem in these papers. However, for this paper we have hundreds of such mergers with small $q$, therefore a simple workaround had to be applied for the traditional treatment. We now use a new expression $A_{\mathrm{MS,3}}$ for the age of the merger product as:
@@
This treatment for the age of the merged body ensures that (i) if $q\rightarrow 1$ (i.e., the two colliding stars have similar masses) we get Eq. 8 (i.e. the age of the merger product is taken from the classical formula) , while (ii) for $q\rightarrow 0$ (i.e., the primary star is much more massive than the secondary) the age of the VMS remains nearly unchanged. Note that Mapelli (2016) used a different rejuvenation algorithm based on earlier papers than Hurley et al. (2005); Glebbeek et al. (2008), which is less general than our algorithms.
Unlike the original treatment in the direct $N$-body simulations using Nbody4, as presented in Hurley et al. (2005), we incorporate in mass loss for hyperbolic collisions, so collisions where the sum of the binding energy of the binary and orbital kinetic energy are positive. If we have a hyperbolic collision between two MS stars, then the code reduces the gained mass by the primary depending on the total energy in the two body center of mass and the binding energy of the secondary, resulting in:
@@ Appendix D Comparison of stellar radii old and upgraded
In SSE the radius of a star at the end of the main sequence is given by the function $R_{\rm TMS}$ (Hurley et al. 2000a, Eqs. 9a and 9b) which can reach very high values in the VMS range: $2\!\cdot\!10^{7}\,\mathrm{R}_{\odot}$ for 50,000 $\mathrm{M}_{\odot}$ at $Z=0.01$. We added an extra function to flatten this out for high masses ($M>3,000\,\mathrm{M}_{\odot}$) where:
@@
This ensures that $R_{\rm TMS}$ cannot become larger than a maximum of $2\cdot 10^{5}\,\mathrm{R}_{\odot}$ which is reached for 50,000 $\mathrm{M}_{\odot}$. This roughly coincides with the radii of VMS models with $M/{\rm M}_{\odot}\geq 10^{4}$, at in the beginning of the Kelvin-Helmholtz contraction phase (Fricke 1973; Hara 1978). (cf. Fig. 9). It was also necessary to impose that the coefficient $a_{24}$ in Eq. 9b of Hurley et al. (2000a) could not become negative (which could previously occur at low $Z$ and cause problems with the radius calculation for VMSs). The term $a_{24}$ is an interpolation term for the $R_{\rm TMS}$. The function to describe the evolution of the radius on the MS is quite complicated (Hurley et al. 2000a, Eq. 13 and associated equations), involving a number of higher order terms and a component aimed at producing the MS hook feature. All of which has the potential to create inconsistencies for very high masses so we simplified the treatment for $M>100\,\mathrm{M}_{\odot}$ by phasing out the hook component (allowing $\Delta R$ to go smoothly to zero across the range $120100<M/\mathrm{M}_{\odot}<$ 120), allowing the $\beta$ term to go to zero (across the same mass range) and capping the $\alpha$ term at its $100\,\mathrm{M}_{\odot}$ value. The $\gamma$ term in Eq. 13 of Hurley et al. (2000a) is already zero in this high mass range. Fig. 9 illustrates the improvements for $M/{\rm M}_{\odot}>500$, showing that we now get a steady increase of stellar radius with time up to the limit defined by Eq. 13. Note that the left panel in Fig. 9 is solely for illustrative purposes, to show what happens if standard SSE would be used for such high masses, for which it has been never intended.
@@ Appendix E Comparison of ageing formulae
Note that Mapelli (2016) use a rejuvenation algorithm based on earlier papers, in which $X=A_{\mathrm{MS,1}}/\tau_{\mathrm{MS,1}}$ only (no dependence on secondary parameters), and a factor 1.0 instead of 0.1 in Eq. 16 was used. Our new model of rejuvenation Eq. 10 becomes
@@
respectively. Now, in Fig.10 $f_{\mathrm{original}}(q,y)$ and $f_{\mathrm{updated}}(q,y)$ are plotted for three values of $y$, i.e. $y\in(-1,-2,-3)$. We can clearly see that $f_{\mathrm{updated}}(q,y)$ produces much better results as predicted for $q=0$ and $q=1$ above. We recommend this correction from Eq. 10 or similar corrections to the original Eq. 8 for everyone still using the original treatment published in Hurley et al. (2002a, 2005); Glebbeek et al. (2008).

### v2

@@ A.1 Treatment of stellar winds for He star
The formation of VMS and posterior collapse as BH, is highly influenced by stellar winds, when it becomes an He star, which vary depending on the star metallicity. We updated the old treatment used in N-body for stellar wind from Hamann1998; Vink2005 with the recipe from Sander2020,
@@ Appendix B Upgrade of stellar radius evolution
Stellar evolution parameters, including the stellar radius, are provided by the SSE algorithm (Hurley2000) with updates as described in Kamlah2022a. This algorithm was fitted to the results of detailed stellar models of up to 50.0 $\mathrm{M}_{\odot}$ and tested for reliable behaviour up to 100.0 $\mathrm{M}_{\odot}$. Thus, using SSE beyond this mass range was historically at the peril of the user and inconsistencies in the stellar radius for masses well above 100.0 $\mathrm{M}_{\odot}$ had been noted previously (e.g. Gieles, private correspondence). For the purpose of the Nbody6++GPU models reported in this work, we therefore conducted a series of tests of SSE in the VMS regime and implemented some minor additions to SSE to safeguard the routines when used for VMS evolution. In Appendix D, including Fig. 9, we explain in some more detail how the upgrade has been done. Note that this is aimed primarily at making the routines numerically safe to use as well as removing any obvious inconsistencies. In the long term, we plan to improve the code with routines for very low mlicities (Tanikawa2020) and massive star formation.
@@ Appendix C Upgrade of rejuvenation via stellar collisions
We use the BSE package by Hurley2002a; Hurley2005, with upgrades related to the treatment of rejuvenation of main sequence stars upon collision with other MS stars and mass loss during hyperbolic collisions, as will be described in this Appendix. Rejuvenation originally was intended to describe the formation of blue straggler stars after stellar collisions (Hurley2005). These stars lie above a stellar populations main sequence turn-off point (see e.g., the observational evidence from Giesers2018). This observation implies that the effective age, $A_{\mathrm{MS}}$, of these MS stars after a collision and a merger, is younger than the average age of the stellar population. This concept (described in more detail below) delivers unphysical results for collisions of a VMS with small mass MS stars, i.e. those where the mass ratio $q$,
@@
becomes very small ($q\ll 1$). In the following, we will analyze in more detail the traditional rejuvenation treatment (Eq. 8, see also discussion in Hurley2005), explain why it fails for very small $q$, and describe our workaround (Eq. 10).
@@
    MS stars with convective cores $\left(M>1.25\penalty 10000\ \mathrm{M}_{\odot}\right)$: upon collision and subsequent mass gain by the primary MS star, its core grows in size. The latter leads to enhanced convective mixing of unburnt hydrogen fuel from the outer stellar layers to the core. This process will lead to the rejuvenation of the star. BSE numerically translate this rejuvenation by conserving the amount of burnt hydrogen and grown the stellar core of the primary star. The new age value is thus directly proportionally to the remaining fraction of unburnt hydrogen at the centre.
@@
    fully convective MS stars $\left(M<0.35\penalty 10000\ \mathrm{M}_{\odot}\right)$: these low-mass stars are treated in the same way as the stars with convective cores.
If we work under the base assumption that hydrogen in the MS core is uniformly distributed after a collision, which in turn is a result of the assumption of instantaneous (convective) mixing of the hydrogen fuel from envelope to core, the BSE code terminates the MS phase of the star and enters the Hertzsprung gap (HG) phase after 10 % of the total amount of hydrogen in the MS star have been burnt (Hurley2000; Hurley2005).
The computational application is described in the following. We first consider the collision of two MS stars with (effective) ages $A_{\mathrm{MS,1}}$ and $A_{\mathrm{MS,2}}$, MS life-times $\tau_{\mathrm{MS,1}}$ and $\tau_{\mathrm{MS,2}}$, as well as masses $M_{\mathrm{MS,1}}$ and $M_{\mathrm{MS,2}}$, respectively. Here, the subscripts $1$ and $2$ denote the primary (and thus more massive) MS star and the secondary, respectively. The collision produces the ”third” MS star with effective age $A_{\mathrm{MS,3}}$, MS life-time $\tau_{\mathrm{MS,3}}$ and mass $M_{\mathrm{MS,3}}$. If we know these parameters, (like suggested in Hurley2005) we can evaluate the final effective age $A_{\mathrm{H,MS,3}}$ of the produced MS star with the following equation (see also Glebbeek2008, for more information):
@@
Such nonphysical age reduction in the previous rejuvenation process has another unwelcome consequence, where the stellar radius of the star is always reset to a value that corresponds to the star’s effective age. However, a MS star grows in size as it ages and approaches the Hertzsprung gap (see e.g. Pols1998; Hurley2000; Tanikawa2020). The stellar size increment accelerates the closer the star is located to the Hertzsprung gap. Therefore, the star will increase in size measured by the stellar radius in several orders of magnitude for VMSs. This implies that a MS star produced via repeated stellar collisions, whose effective age is treated with Eq. 8, will be much too small, which in turn greatly also affects its scattering cross-section and therefore, its mass growth, in dense stellar environment.
In our previous direct $N$-body works, where only a limited number of stellar collisions have been produced, no MS star merger chains have been produced that go beyond 10 compared with the densest star cluster model presented here with $r_{\mathrm{h,i}}=0.1\penalty 10000\ (\mathrm{pc})$ collisions (see e.g. Rizzuto2021; Rizzuto2022; Kamlah2022a; Kamlah2022b; ArcaSedda2023; ArcaSedda2024; ArcaSedda2024a, for recent direct $N$-body work, which includes comparable astrophysical complexity as the work presented here, i.e. full stellar evolution of single and binary stars, tidal fields, mass spectrum and so on.). Therefore the aforementioned rejuvenation issue was not a problem in these papers. However, for this paper we have hundreds of such mergers with small $q$, therefore a simple workaround had to be applied for the traditional treatment. We now use a new expression $A_{\mathrm{MS,3}}$ for the age of the merger product as:
@@
This treatment for the age of the merged body ensures that (i) if $q\rightarrow 1$ (i.e., the two colliding stars have similar masses) we get Eq. 8 (i.e. the age of the merger product is taken from the classical formula) , while (ii) for $q\rightarrow 0$ (i.e., the primary star is much more massive than the secondary) the age of the VMS remains nearly unchanged. Note that Mapelli2016 used a different rejuvenation algorithm based on earlier papers than Hurley2005; Glebbeek2008, which is less general than our algorithms.
Unlike the original treatment in the direct $N$-body simulations using Nbody4, as presented in Hurley2005, we incorporate in mass loss for hyperbolic collisions, so collisions where the sum of the binding energy of the binary and orbital kinetic energy are positive. If we have a hyperbolic collision between two MS stars, then the code reduces the gained mass by the primary depending on the total energy in the two body center of mass and the binding energy of the secondary, resulting in:
@@ Appendix D Comparison of stellar radii old and upgraded
In SSE the radius of a star at the end of the main sequence is given by the function $R_{\rm TMS}$ (Hurley2000, Eqs. 9a and 9b) which can reach very high values in the VMS range: $2\!\cdot\!10^{7}\,\mathrm{R}_{\odot}$ for 50,000 $\mathrm{M}_{\odot}$ at $Z=0.01$. We added an extra function to flatten this out for high masses ($M>3,000\,\mathrm{M}_{\odot}$) where:
@@
This ensures that $R_{\rm TMS}$ cannot become larger than a maximum of $2\cdot 10^{5}\,\mathrm{R}_{\odot}$ which is reached for 50,000 $\mathrm{M}_{\odot}$. This roughly coincides with the radii of VMS models with $M/{\rm M}_{\odot}\geq 10^{4}$, at in the beginning of the Kelvin-Helmholtz contraction phase (Fricke1973; Hara1978). (cf. Fig. 9). It was also necessary to impose that the coefficient $a_{24}$ in Eq. 9b of Hurley2000 could not become negative (which could previously occur at low $Z$ and cause problems with the radius calculation for VMSs). The term $a_{24}$ is an interpolation term for the $R_{\rm TMS}$. The function to describe the evolution of the radius on the MS is quite complicated (Hurley2000, Eq. 13 and associated equations), involving a number of higher order terms and a component aimed at producing the MS hook feature. All of which has the potential to create inconsistencies for very high masses so we simplified the treatment for $M>100\,\mathrm{M}_{\odot}$ by phasing out the hook component (allowing $\Delta R$ to go smoothly to zero across the range $100<M/\mathrm{M}_{\odot}<120$), allowing the $\beta$ term to go to zero (across the same mass range) and capping the $\alpha$ term at its $100\,\mathrm{M}_{\odot}$ value. The $\gamma$ term in Eq. 13 of Hurley2000 is already zero in this high mass range. Fig. 9 illustrates the improvements for $M/{\rm M}_{\odot}>500$, showing that we now get a steady increase of stellar radius with time up to the limit defined by Eq. 13. Note that the left panel in Fig. 9 is solely for illustrative purposes, to show what happens if standard SSE would be used for such high masses, for which it has been never intended.
@@ Appendix E Comparison of ageing formulae
Note that Mapelli2016 use a rejuvenation algorithm based on earlier papers, in which $X=A_{\mathrm{MS,1}}/\tau_{\mathrm{MS,1}}$ only (no dependence on secondary parameters), and a factor 1.0 instead of 0.1 in Eq. 16 was used. Our new model of rejuvenation Eq. 10 becomes
@@
respectively. Now, in Fig.10 $f_{\mathrm{original}}(q,y)$ and $f_{\mathrm{updated}}(q,y)$ are plotted for three values of $y$, i.e. $y\in(-1,-2,-3)$. We can clearly see that $f_{\mathrm{updated}}(q,y)$ produces much better results as predicted for $q=0$ and $q=1$ above. We recommend this correction from Eq. 10 or similar corrections to the original Eq. 8 for everyone still using the original treatment published in Hurley2002a; Hurley2005; Glebbeek2008.
@@
Appendix F Hardware settings
The simulations are executed on the Raven high performance computing system at the Max Planck Computing and Data Facility (MPCDF) [^3] Raven: https://www.mpcdf.mpg.de/services/supercomputing/raven. Here, we use from one to three full GPU computing node for the simulation, which gives the optimal performance and most cost-effective use of the computing resources. Each GPU computing node is equipped 4 Nvidia A100 GPUs (4 × 40 GB HBM2 memory per node and NVLink). Furthermore, each node has one CPU hosts, which is an Intel Xeon IceLake Platinum 8360Y with 72 CPU cores. The RAM is 512 GB per node or 256 GB of RAM depending on the node you get assigned to via the slurm system. For a detailed account including benchmarking concerning the optimal performance of the Nbody6++GPU code on hybrid, GPU-accelerated supercomputers, please refer to the recent review by Spurzem2023 and sources therein. Even using this state-of-the-art high performance computing (HPC) infrastructure, the simulation take up to 2 months of real time to reach a star cluster age of 5 Myr. Tests and developments on machines such as juwels-booster in Germany  [^4] Juwels-booster: https://apps.fz-juelich.de/jsc/hps/juwels/booster-overview.html, Lumi in Finland and Leonardo in Italy were done. We expect more data from these infrastructures soon.